OpenStudy (anonymous):

help me!! please! what is the solution set of -3ln6-1ln3 divided by 2ln3-ln6

OpenStudy (anonymous):

you must first use the rule alnb = ln b^a then you use the rule lna - lnb = ln a/b and then put it in a calculator. So this is how it goes (ln 6^-3) - (ln3^1) = ln(6^-3)/3(^1) = ln(6^-3)/3 (ln3^2)-ln6) =(ln9/6) = ln 3/2 then divide the first expression by the second one by using your calculator.